Daily Enterprises is purchasing a $ 9.9 million machine. It will cost $51,000 to transport and install the machine. The machine has a depreciable life of five years and will have no salvage value. The machine will generate incremental revenues of $4.1 million per year along with incremental costs of $1.1 million per year. If Daily's marginal tax rate is 35 %, what are the incremental earnings (net income) associated with the new machine?

AIE = (Revenues - Costs - Depreciation) x (1 - tax rate)
Answer #1

Cost of Machine = Purchase price + Cost of Installing = 9,900,000 + 51000 = $9,951,000

Depreciation per year = 9,951,000 / 5 = $1,990,200

Incremental Cash inflows from year 1-5 are:

Incremental Revenues = $4,100,000

Less: Incremental Costs = $1,100,000

Less: Depreciation = $1,990,200

Profit before taxes = $1,009,800

Less: Taxes @35% = $353,430

Profit after taxes = $656,370

Add: Depreciation = $1,990,200

Incremental Cash Inflows = $2,646,570